OpenStudy (anonymous):

Simplify and write in radical form: ((x^3 * y^-2) / xy)^-(1/5)

OpenStudy (anonymous):

\[(x ^{3}y ^{-2}/xy)^{-1/5}\] this equation?

OpenStudy (anonymous):

Yes, that's the one.

OpenStudy (anonymous):

Take it one step at a time:\[(\frac{x^{3}}{y^{2}}\frac{1}{xy})^{\frac{-1}{5}}\]\[(\frac{x^{3}}{xy^{3}})^{\frac{-1}{5}}\]\[(\frac{x^{2}}{y^{3}})^{\frac{-1}{5}}\]\[(\frac{y^{3}}{x^{2}})^{\frac{1}{5}}\]\[\sqrt[5]{\frac{y^{3}}{x^{2}}}\]